Wendy and her mother are planning for the annual family reunion. there will be 84 people attending.hot dogs are sold in packages of 12 and buns are sold in bags of 8.how many bags of buns should Wendy's mom buy to serve 4 packages of hot dogs?
1
Expert's answer
2012-10-29T11:40:03-0400
There are

H = 12*4 = 48

hot dogs in a 4 packages. Therefore, she needs

B = 48/8 = 6

bags of buns to serve 4 packages of hot dogs.